OpenStudy (anonymous):

if A^4=0, then I-A is ivertible, (I-A)^-1=I+A+A^2+A^3. T or F, prove it

OpenStudy (anonymous):

Let me write it clearly then you can continue with him or her: \[(I-A)^{-1}=I+A+A^2+A^3\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Oh, so finally you fell down here.. :P

OpenStudy (anonymous):

i think it is T

OpenStudy (anonymous):

but can not prove it

OpenStudy (anonymous):

How you can say that?? Because of your positive and optimistic nature towards life?? :P

OpenStudy (ikram002p):

hmm is this a given or need to prove ? (I-A)^-1=I+A+A^2+A^3

OpenStudy (anonymous):

The question is asking whether RHS is equal to LHS or not..

OpenStudy (anonymous):

We have to solve for \((1-A^{-1})\) and get the right hand side thing. Or you can do reverse also.. :)

OpenStudy (anonymous):

But how can we do this, I don't know still.. :P

OpenStudy (anonymous):

Oh, -1 is outside brackets..

OpenStudy (anonymous):

it is quite equal to I^4-A^4

OpenStudy (ikram002p):

how ? (I-A)^-1 = (I-A^-1)

OpenStudy (anonymous):

I said -1 is outside brackets.. :P Do you wear glasses, ikram?? :P

OpenStudy (anonymous):

:)) two guys...

OpenStudy (ikram002p):

:3 nvm

OpenStudy (anonymous):

Never mind from my side too, I was just kidding.. :)

OpenStudy (ikram002p):

@wonnguyen1995 and ur 3 bhahaha

OpenStudy (anonymous):

May be, we need experts' advice here.. He he he.. :) @ganeshie8

OpenStudy (anonymous):

Which topic are you studying @wonnguyen1995

OpenStudy (anonymous):

Apart from this I mean.. Have you studied of Binomial Theorems??

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

This topic is under Matrices And Determinants, right??

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

but my major just need basic for this

OpenStudy (anonymous):

You are learning Matrices And Determinants and Binomial has not been taught to you yet??

OpenStudy (anonymous):

i have not studied determinant and BO

OpenStudy (anonymous):

I have a strong feeling that Binomial expansion has a significance here.. :)

OpenStudy (anonymous):

you know any guy who could solve it here?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Let see the expansion of \((1-x)^{-1}\).. \[(1-x)^{-1} = 1 + x + x^2 + x^3 + x^4 + x^5 + ........\]

OpenStudy (anonymous):

This is matching to your right hand side.. Miracle..!!! :P

OpenStudy (anonymous):

If you can replace x with A, and you know \(A^4 = 0\), so every term afer \(A^4\) will be 0 I think..

OpenStudy (anonymous):

So, you will remain upto exponent of 3 only.. \[(1-A)^{-1} = 1 + A + A^2 + A^3\]

OpenStudy (anonymous):

\[I^4-A^4 = (I-A)(I^3+I^2*A+I*A^2+A^3))\]

OpenStudy (anonymous):

Really??

OpenStudy (anonymous):

Wait, let me see it: \[I^4 - A^4 = (I-A)(I+A)(I^2 + A^2) \implies \color{green}{I^4 - A^4 = (I-A)(I + A + A^2 + A^3)}\]

OpenStudy (anonymous):

Yeah you went really great.. Appreciable..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Now as \(A^4 = 0\), and divide both sides by \((I-A)\) as it is invertible so you can divide.. Just fantastic approach.. :)

OpenStudy (anonymous):

Got?? Or still in doubt??

OpenStudy (anonymous):

:) you are enthusiastic

OpenStudy (anonymous):

Me ?? :P